# Fast Point Cloud to Mesh Reconstruction for Deformable Object Tracking

### v2

@@ IV-B3 Comparison of Different Works
In Table I, the factors of generalizability and speed of the most relevant works to ours are compared. As discussed before, these two factors are essential for robotics applications. If a method works slower than then its application may not be considered in a robotics context. Furthermore, the generalizability of a method to different object categories is essential for robotics manipulation tasks where any object of an unpredictable category could need manipulation. Therefore, methods that have been designed for one single object category are not suitable for the problem addressed in our work. Some of the works mentioned in Table I, address our problem only in part. For instance, in the works, OccFlow [14], CaDeX [12] and Topology[18], the template mesh is learned in addition to the deformation and hence every trained model works on only one category of objects. For instance, they have individually trained models for each of the categories humans, animals, doors, etc. In contrast, we consider the template mesh learning as a separate problem resolved by SOTA methods such as Shape As Points [17] and therefore our model remains generalizable to different object categories. Furthermore, despite both our method and CaDeX using Real-NVP [6], separating template learning from deformation learning in our approach leads to an increased speed over CaDeX. Moreover, the work, Point2Mesh [8], does not do any vertex tracking in addition to being slow for robotic applications. The NMF [7] is also not time efficient as it works at around which makes it incompatible with robotics settings.
Nevertheless, in order to have a baseline to compare against in terms of quality, we retrained the Occupancy flow [14] model from scratch on our dataset D. In contrast to their training dataset, we have different object categories within the dataset D. The Occupancy flow [14] is the most recent work in the 4D reconstruction SOTA to the best of our knowledge. In Table V, our work is quantitatively compared against Occupancy flow [14]. Our work has a lower chamfer loss. In Fig. 16 of the Appendix, our deformed meshes are qualitatively compared against the deformed meshes generated by Occupancy flow [14] for two YCB objects. As illustrated, the bleach cleanser and the scissors are, respectively, missing details and partially reconstructed. Furthermore, the model does not generalize to the 18th-frame deformation of the bleach cleanser. We speculate that because Occupancy flow [14] needs to learn the template mesh, its generalizability capacity to different YCB objects is limited.
@@ Appendix A Prior method results issues
In this section, some visualizations are provided in order to give a clearer understanding of the different method performances. In deforming methods, the mesh is generated by moving around the vertices of the initial mesh in order to fit the target point cloud. For instance, for reconstructing a car point cloud, the deformation network starts off with a car mesh as can be seen in Fig. 8a or for reconstructing a chair point cloud, the network starts off with a similar chair mesh as can be seen in Fig. 8b. The method Point2mesh[8] results in non-water tight meshes as can be seen in Fig. 9. The NMF[7] method, which deforms a sphere, can not reconstruct meshes of a higher genus than zero as can be seen in Fig. 10. The NMF[7] is incapable to fit thin structures such as lamps as can be seen in Fig. 11. The neural parts[16] method deforms five template spheres as can be seen in Fig. 12. The latter gives the possibility of reconstructing meshes of more complex topology as can be seen in Fig. 13
@@
In Fig. 16 the results of our method and the results of the method Occupancy flow[14] both trained on our dataset D are visualized.

### v3

@@ IV-B3 Comparison of Different Works
In Table I, the factors of generalizability and speed of the most relevant works to ours are compared. As discussed before, these two factors are essential for robotics applications. If a method works slower than then its application may not be considered in a robotics context. Furthermore, the generalizability of a method to different object categories is essential for robotics manipulation tasks where any object of an unpredictable category could need manipulation. Therefore, methods that have been designed for one single object category are not suitable for the problem addressed in our work. Some of the works mentioned in Table I, address our problem only in part. For instance, in the works, OccFlow [occflow], CaDeX [cadex] and Topology[topologyPreservation], the template mesh is learned in addition to the deformation and hence every trained model works on only one category of objects. For instance, they have individually trained models for each of the categories humans, animals, doors, etc. In contrast, we consider the template mesh learning as a separate problem resolved by SOTA methods such as Shape As Points [shapeaspoints] and therefore our model remains generalizable to different object categories. Furthermore, despite both our method and CaDeX using Real-NVP [realnvp], separating template learning from deformation learning in our approach leads to an increased speed over CaDeX. Moreover, the work, Point2Mesh [point2mesh], does not do any vertex tracking in addition to being slow for robotic applications. The NMF [nmf] is also not time efficient as it works at around which makes it incompatible with robotics settings.
Nevertheless, in order to have a baseline to compare against in terms of quality, we retrained the Occupancy flow [occflow] model from scratch on our dataset D. In contrast to their training dataset, we have different object categories within the dataset D. The Occupancy flow [occflow] is the most recent work in the 4D reconstruction SOTA to the best of our knowledge. In Table V, our work is quantitatively compared against Occupancy flow [occflow]. Our work has a lower chamfer loss. In Fig. 16 of the Appendix, our deformed meshes are qualitatively compared against the deformed meshes generated by Occupancy flow [occflow] for two YCB objects. As illustrated, the bleach cleanser and the scissors are, respectively, missing details and partially reconstructed. Furthermore, the model does not generalize to the 18th-frame deformation of the bleach cleanser. We speculate that because Occupancy flow [occflow] needs to learn the template mesh, its generalizability capacity to different YCB objects is limited.
@@ V Conclusion and Future Work
VI Acknowledgment
This study was supported by RobotX Research Grant, RX-03-22, “LPDR - Learning Physics-based Dense Representations for Robotic Manipulation of Soft and Articulated Objects”.
[title=Reference]
@@ Appendix A Prior method results issues
In this section, some visualizations are provided in order to give a clearer understanding of the different method performances. In deforming methods, the mesh is generated by moving around the vertices of the initial mesh in order to fit the target point cloud. For instance, for reconstructing a car point cloud, the deformation network starts off with a car mesh as can be seen in Fig. 8a or for reconstructing a chair point cloud, the network starts off with a similar chair mesh as can be seen in Fig. 8b. The method Point2mesh[point2mesh] results in non-water tight meshes as can be seen in Fig. 9. The NMF[nmf] method, which deforms a sphere, can not reconstruct meshes of a higher genus than zero as can be seen in Fig. 10. The NMF[nmf] is incapable to fit thin structures such as lamps as can be seen in Fig. 11. The neural parts[neuralparts] method deforms five template spheres as can be seen in Fig. 12. The latter gives the possibility of reconstructing meshes of more complex topology as can be seen in Fig. 13
@@
In Fig. 16 the results of our method and the results of the method Occupancy flow[occflow] both trained on our dataset D are visualized.